OpenStudy (anonymous):

Hal invests $15,000.00 in a savings account which pays 7% compounded continuously. Consider the following formula, where A is the ending account balance after t years, P is the initial amount of money invested, and r is the interest rate. A = P(2.71)rt How much money would he have in his savings account after 2 years? $17,246.74 $17,948.40 $15,610.26 $16,084.19

OpenStudy (solomonzelman):

It's supposed to be to the power of rt , correct ?

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

Yeahh.

OpenStudy (wolf1728):

Continuously Compounded Rate of 7 % = (e^r) -1 (2.718281828^.07)-1 =0.0725081812 or 7.25081812 % annual rate A = 15,000 * (1.0725081812)^2 A = 15,000 * (1.1502737987) A = 17,254.11 but that isn't one of the choices is it?

OpenStudy (anonymous):

nope its not,

OpenStudy (anonymous):

i figured it out. thanks for helping.

OpenStudy (wolf1728):

Wow - okay - the continuously compounded interest too? All right

OpenStudy (wolf1728):

I got the answer. I'm surprised that they rounded the value of 'e' as 2.71 I computed the compounded interest as (2.718281828)^.07 which = 1.1497825171 My method (using a more precise value of 'e') comes to 1.1502737988 A small difference. Well it amounts to a 2 year difference of $7.37